Given the following equation:            8 Fe + S8 ---> 8 FeS

What mass of Fe is needed to react with 16.0 grams of sulfur (S8)?

(Fe= 55.8, S=32)

 


1
Expert's answer
2021-10-05T01:36:33-0400

M(Fe) = 55.8 g/mol

M(S8) = 8 × M(S) = 8 × 32 = 256 g/mol


Solution:

Balanced chemical equation:

8Fe + S8 ⟶ 8FeS


According to stoichiometry:

(16.0 g S8) × (1 mol S8 / 256 g S8) × (8 mol Fe / 1 mol S8) × (55.8 g Fe / 1 mol Fe) = 27.9 g Fe


Answer: 27.9 grams of Fe is needed to react with 16.0 grams of S8